# Conformal Prediction Sets Improve Human Decision Making

### v1

@@ A.2 Conformal Prediction Method and Hyperparameters
To generate conformal prediction sets we used the well-known RAPS procedure (Angelopoulos et al., 2021), which provably produces sets of lower average size than the top-$k$ method for an equivalent coverage level. RAPS has three hyperparameters: a temperature $T$ used for scaling the model’s logits before applying the softmax; a ranking regularizer $k_{\mathrm{reg}}$, and a regularization term weight $\lambda$.
@@
Table 5: Hyperparameter Settings, and Empirical Coverage
@@ Appendix B Human Subject Experiments
In this section we provide additional details on our human subject experiments to complement the descriptions in Sections 3 and 4.
@@
When running any experiment involving human subjects, ethical considerations are of utmost importance. We considered research ethics through the experimental process, and followed the ethics guidelines for experiments involving humans published by the NeurIPS 2023 conference organizers. In particular, we followed all existing protocols at our institution for such research. Although our institution does not have an internal review board (IRB) process, in its place we took the following steps: understanding the existing process in place at our institution, ensuring that our datasets only contained content appropriate for showing participants, piloting the tests ourselves, informing participants about what data would be collected and what data they would be shown, collecting consent from participants, paying participants a fair wage, and evaluating the demographic distribution of participants in our experiments to be aware of potential biases.
@@
We recruited participants through Prolific, which produces the highest quality data according to scientific comparisons of behavioural data collection platforms (Eyal et al., 2021; Douglas et al., 2023). We did not filter the population accepted into our study, other than that we required fluency in English, the language we used for task instructions. When participants voluntarily agreed to join our study, they were randomly assigned to one of three tasks, and one of three treatments. No participant was involved in more than one test. In total, we recruited and paid 600 participants. Compensation followed the guidelines enforced by Prolific. Participants were promised a flat amount of pay based on the median completion time of the test, and were also offered bonus pay proportional to the number of correct answers they gave as a financial incentive to perform well on the tasks. Overall, the average rate of pay was 7.80 GBP/hr.
All 600 participants consented to the collection and dissemination-in-aggregate of their demographic data. As seen in Table 6, our study population was approximately balanced in terms of gender, and covered a wide range of ages. Since all the tasks we designed rely only on general human knowledge (and fluency in English), we do not expect any effect of age or gender on task performance (here measured by accuracy). To verify this null hypothesis, we compare the normalized accuracy for each group in Table 6. Because each task and treatment may have a different inherent difficulty, we normalize the accuracy within each task/treatment cohort before aggregating over demographic groups. We show the mean and standard deviation, where a value of 1 indicates the population performed 1 standard deviation above the mean of their cohort. Unsurprisingly, the results are consistent with there being no effect of gender or age on task performance.
@@
Participants were required to use a desktop or laptop computer to complete our experiment (not a mobile device or tablet). To begin, participants were shown a statement on the data we were collecting, including information on how we planned to store, disseminate, and release that data publicly for scientific purposes. Participants had the option to consent to this treatment, or remove themselves from the study. Then, instructions were shown introducing the task, and we provided labelled examples of each class. Next, participants went through 20 practice trials of the same format as the actual test so they could further learn about the dataset and classes. Example and practice stimuli were taken from $\mathcal{D}_{\mathrm{test}}$, were the same for all participants, and were never reselected as a test trial. Finally, the test trials were conducted using randomized samples from $\mathcal{D}_{\mathrm{test}}$; for both GoEmotions and Few-NERD $m=50$ trials were used, while for ObjectNet we assigned $m=100$ trials because they could be completed more quickly. Our experimental data is based only on the test trials. Figure 12 shows the screens that were displayed to participants in the Few-NERD experiment; the other two experiments followed the same template, and examples of their main trial screens are given in Figure 2 and Figure 13.
@@
Each trial showed the participant a stimulus $x$ and all possible classes $\mathcal{Y}$, forcing the participant to choose one class with unlimited time to make their choice. Participants were given the correct answer after they responded, both during the practice and testing phase. Training of this kind has been shown to have a large positive effect on the quality of data collected (Mitra et al., 2015). It is possible that the participant may learn about the task while taking the test and change the way they answer to maximize accuracy. We see this as a desirable effect; ultimately we want participants to perform as well as possible. Additional justifications for providing correct answers during the testing phase are given in (Stein et al., 2023).
@@ Appendix C Additional Analysis
Figure 3 summarized our main experimental results regarding human accuracy on our tasks. In Figure 14 we show the histograms of $N=50$ accuracy observations for each task, making pairwise comparisons between treatments. According to the significance tests compiled in Table 2, the conformal treatment shows statistically significant improvement in accuracy compared to the other treatments on each of the tasks. From the histograms we also see that the distributions of observations are visually consistent with binomial distributions, as discussed in Section 3.
Figure 14: Histograms of accuracy observations across three treatments for each task. Vertical lines indicate the mean of each distribution.
Our results on response times were displayed in Figure 4. Once again we show the histograms of $N=50$ response time observations for each task in Figure 15. Unlike for accuracy, we do not see strong trends for prediction sets increasing or decreasing decision speed across all cases, and in most cases we do not see significant differences between the means of the distributions (Table 3).
Figure 15: Histograms of response time observations across three treatments for each task. Vertical lines indicate the mean of each distribution.
In our pre-registration at osf.io/fkdhv we stated that data analysis would be conducted using the one-way ANOVA method. In the ensuing time since pre-registration, and based on the discussion in Section 3 we found that applying Welch’s $t$-test was a suitable, and simpler statistical approach, and therefore used it for our final analysis instead.

### v3

@@ A.2 Conformal Prediction Method and Hyperparameters
To generate conformal prediction sets we used the well-known RAPS procedure (Angelopoulos et al., 2021), which provably produces sets of lower average size than the top-$k$ method for an equivalent coverage level. RAPS has three hyperparameters: a temperature $T$ used for scaling the model’s logits before applying the softmax; a set size regularizer $k_{\mathrm{reg}}$; and a regularization weight $\lambda$.
@@
Table 6: Hyperparameter Settings and Empirical Coverage
@@ Appendix B Human Subject Experiments
In this section we provide additional details on our human subject experiments to complement the descriptions in Sections 3 and 4. We have released our aggregated experimental data in this GitHub repository.
@@
When running any experiment involving human subjects, ethical considerations are of utmost importance. We considered research ethics throughout the experimental process, and followed the ethics guidelines for experiments involving humans published by the NeurIPS 2023 conference organizers, which accords with the standards of ICML 2024. In particular, we followed all existing protocols at our institution for such research. Although our institution does not have an internal review board (IRB) process, in its place we took the following steps: understanding the existing process in place at our institution, ensuring that our datasets only contained content appropriate for showing participants, piloting the tests ourselves, informing participants about what data would be collected and what data they would be shown, collecting consent from participants, paying participants a fair wage, and evaluating the demographic distribution of participants in our experiments to be aware of potential biases.
@@
We recruited participants through Prolific, which produces the highest quality data according to scientific comparisons of behavioural data collection platforms (Eyal et al., 2021; Douglas et al., 2023). We did not filter the population accepted into our study, other than that we required fluency in English, the language we used for task instructions. When participants voluntarily agreed to join our study, they were randomly assigned to one of three tasks, and one of three treatments. No participant was involved in more than one test. In total, we recruited and paid 600 participants. Compensation followed the guidelines enforced by Prolific. Participants were promised a flat amount of pay based on the median completion time of the test, and were also offered bonus pay proportional to the number of correct answers they gave as a financial incentive to perform well on the tasks. Overall, the average rate of pay was 7.80 GBP/hr, and 1500 GBP was spent on participant compensation in total, exclusive of fees.
All 600 participants consented to the collection and dissemination-in-aggregate of their demographic data. As seen in Table 7, our study population was approximately balanced in terms of gender (58% Male, 42% Female), and covered a wide range of ages. Since all the tasks we designed relied only on general human knowledge (and fluency in English), we do not expect any effect of age or gender on task performance (here measured by accuracy). To verify this null hypothesis, we compare the normalized accuracy for each group in Table 7. Because each task and treatment may have a different inherent difficulty, we normalize the accuracy within each task/treatment cohort before aggregating over demographic groups. We show the mean and standard deviation, where a value of 1 indicates the population performed 1 standard deviation above the mean of their cohort. Unsurprisingly, the results are consistent with there being no effect of gender or age on task performance.
@@
Participants were required to use a desktop or laptop computer to complete our experiment (not a mobile device or tablet). To begin, participants were shown a statement on the data we were collecting, including information on how we planned to store, disseminate, and release that data publicly for scientific purposes. Participants had the option to consent to this treatment, or remove themselves from the study (see Figure 11). We did not collect Personal Identifiable Information (PII) such as name, address, birth date, governmental identification numbers, or banking information. There were no potential risks of participating that we needed to disclose to participants. Then, instructions were shown introducing the task, and we provided labelled examples of each class. Next, participants went through 20 practice trials of the same format as the actual test so they could further learn about the dataset and classes. Example and practice stimuli were taken from $\mathcal{D}_{\mathrm{test}}$, were the same for all participants, and were never reselected as a test trial. Finally, the test trials were conducted using randomized samples from $\mathcal{D}_{\mathrm{test}}$; for both GoEmotions and Few-NERD $m=50$ trials were used, while for ObjectNet we assigned $m=100$ trials because they could be completed more quickly. Our experimental data is based only on the test trials. Figure 13 shows the screens that were displayed to participants in the Few-NERD experiment; the other two experiments followed the same template, and examples of their main trial screens are given in Figure 2 and Figure 12.
Figure 11: Consent screen shown to participants at the start of the experiment.
@@
Each trial showed the participant a stimulus $x$ and all possible classes $\mathcal{Y}$, forcing the participant to choose one class with unlimited time to make their choice. For ObjectNet, the stimuli were only shown for 0.22 s to increase the difficulty of the task. Participants were given the correct answer after they responded, both during the practice and testing phase. Training of this kind has been shown to have a large positive effect on the quality of data collected (Mitra et al., 2015). It is possible that the participant may learn about the task while taking the test and change the way they answer to maximize accuracy. We see this as a desirable effect; ultimately we want participants to perform as well as possible. Additional justifications for providing correct answers during the testing phase are given in (Stein et al., 2023).
@@ Appendix C Additional Analysis
Figure 3 summarized our main experimental results regarding human accuracy on our tasks. In Figure 14 we show the histograms of $N=50$ accuracy observations for each task, making pairwise comparisons between treatments. According to the significance tests compiled in Table 2, the conformal treatment shows statistically significant improvement in accuracy compared to the other treatments on each of the tasks.
Figure 14: Histograms of accuracy observations across three treatments for each task. Vertical lines indicate the mean of each distribution, $p$ indicates the $p$-value given by Welch’s $t$-test, and $d$ is the effect size given by Cohen’s $d$.
Our results on response times were displayed in Figure 4. Once again we show the histograms of $N=50$ average response time observations for each task in Figure 15. Unlike for accuracy, we do not see strong trends for prediction sets increasing or decreasing decision speed across all cases, and in most cases we do not see significant differences between the means of the distributions (Table 3).
Figure 15: Histograms of response time observations across three treatments for each task. Vertical lines indicate the mean of each distribution, $p$ indicates the $p$-value given by Welch’s $t$-test, and $d$ is the effect size given by Cohen’s $d$.
In our pre-registration we stated that data analysis would be conducted using the one-way ANOVA method. In the ensuing time since pre-registration, and based on the discussion in Section 3 we found that applying Welch’s $t$-test was a suitable, and simpler statistical approach, and therefore used it for our final analysis instead.